OpenStudy (anonymous):

design 5 numbers which have a range of 10, a median of 7 and a mean of 6 please solve

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

ok lets fill in a couple numbers 1 _ 7 _ 11. this gives you your median of 7 and range of 10. now for an average of 6: (1+7+11+a+b)/5=6 a+b=30-19 a+b=11 so we can shoose a=3, b=8 1,3,7,8,11

OpenStudy (anonymous):

or 1,2,7,9,11

OpenStudy (anonymous):

yep that is ok too

OpenStudy (anonymous):

any two a+b=11 with a between 1 and 7, b between 7 and 11

OpenStudy (anonymous):

tks :) it cud b 1,4,7,7,11 also